OpenStudy (anonymous):

y=2x+5 to find the values of y then x=1,2,3 and 4. Write the answers as ordered pairs that would be used to graph the function

OpenStudy (anonymous):

Draw a table with columns for x and y. subst. values and graph the solution

OpenStudy (jack1):

(x, y) if y=2x+5 , and x = 1... y=2(1)+5 y=2+5 y = 7 so ordered pair = (x, y) = (1, 7) in this case now u try when x = 2 @browneyedhtty26

OpenStudy (anonymous):

Graph should look something like this